OpenStudy (anonymous):

\[\sqrt[5]{96}+\sqrt[5]{3}=3\sqrt[5]{3} \] Can someone show me how this answer is correct?

OpenStudy (anonymous):

\[96=3\times32 so \sqrt[5]{96}=\sqrt[5]{3\times32}=\sqrt[5]{3}\times \sqrt[5]{32}=2\sqrt[5]{3}\]

OpenStudy (anonymous):

Thank you so much! Now, I understand. :)